OpenStudy (anonymous):

Choose the correct simplification of the expression −4x2(6x − 5x2 − 5). 20x4 + 24x3 + 20x2 −9x4 + 2x3 − 9x2 20x4 − 24x3 + 20x2 −20x4 + 24x3 − 20x2

OpenStudy (anonymous):

You distribute first.

OpenStudy (anonymous):

What do you get when you distribute?

OpenStudy (anonymous):

is it C?

OpenStudy (anonymous):

This is like say \( -4x^2 \times 6x \ - 4x^2 \times -5x^2 \ -4x^2 \times -5 \)

OpenStudy (anonymous):

I have no clue

OpenStudy (anonymous):

Yes it is C

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

YW